Verify that: -(-x) = x for. (i) x = 11/15 (ii) x = -13/17 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

x = 11/15
We have, x = 11/15
Theadditiveinverseofxis–x (as x+(-x)=0)
Then, the additive inverse of 11/15 is–11/15 (as 11/15 + (-11/15) = 0)
The same equality 11/15 + (-11/15) = 0, shows that the additive inverse of -11/15 is 11/15. Or, - (-11/15) = 11/15
i.e., -(-x) = x

-13/17
We have, x = -13/17
Theadditiveinverseofxis–x (as x+(-x)=0)
Then, the additive inverse of -13/17 is 13/17 (as 11/15 + (-11/15) = 0)
The same equality (-13/17 + 13/17) = 0, shows that the additive inverse of 13/17 is -13/17. Or, - (13/17) = -13/17,
i.e., -(-x) = x

Thanks